OpenStudy (anonymous):

Please assist me solving -4(3 + w) = -27 -w

OpenStudy (anonymous):

-12-4w=-27-w 12+4w=27+w 3w=15 w=5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

how did you get 3

OpenStudy (anonymous):

4w-1w=3w

OpenStudy (anonymous):

and you divided 15 by 3 to get 5?